### v1

# Decoding Cortical Microcircuits: A Generative Model for Latent Space Exploration and Controlled Synthesis

Abstract
A central idea in understanding brains and building artificial intelligence is that structure determines function. Yet, how the brain’s complex structure arises from a limited set of genetic instructions remains a key question. The ultra high-dimensional detail of neural connections vastly exceeds the information storage capacity of genes, suggesting a compact, low-dimensional blueprint must guide brain development. Our motivation is to uncover this blueprint. We introduce a generative model, to learn this underlying representation from detailed connectivity maps of mouse cortical microcircuits. Our model successfully captures the essential structural information of these circuits in a compressed latent space. We found that specific, interpretable directions within this space directly relate to understandable network properties. Building on this, we demonstrate a novel method to controllably generate new, synthetic microcircuits with desired structural features by navigating this latent space. This work offers a new way to investigate the design principles of neural circuits and explore how structure gives rise to function, potentially informing the development of more advanced artificial neural networks.
1 Introduction
The relationship between structure and function is a core idea in both neuroscience and the development of artificial intelligence Gradientsofstructure-functiontetheringacrossneocortex ; NAS1 ; NAS2 . The brain, with its amazing abilities, inspires us to build better AI systems neuronAI1 ; neuronAI2 ; hassabis2017 ; lecun2015 . However, the brain’s structure is incredibly complex complex_connectome1 ; complex_connectome2 . Understanding how all its neurons are connected is a huge task. These maps of connections are called connectomes connectome_def1 ; connectome_def2_original ; connectome_def3 . Studying connectomes is important because they hold clues about how the brain processes information and learns HCP ; USbrain1 .
A key puzzle is how the brain’s intricate wiring instructions are stored. An animal’s genes guide the development of its nervous system Genetic_influences_on_hub_connectivity_of_the_human_connectome ; Genetic_architecture_of_the_structural_connectome . However, the amount of information genes can hold is much smaller than what would be needed to explicitly list every single connection in a fully formed brain info_capacity1 ; info_capacity2 . This observation, known as the "genetic bottleneck" pnas_genetic_btnk , suggests that there must be a simpler, more compact set of rules or a low-dimensional representation that guides how brain networks grow and organize themselves.
This paper aims to uncover such a low-dimensional code for brain circuitry. We focus on cortical microcircuits, which are small, repeating patterns of cortical connections that can be thought of as fundamental building blocks of the brain. To do this, we use data from the MICrONS program microns , which has produced a large and detailed connectome dataset from the visual cortex of a mouse. This dataset provides an unprecedented opportunity to study the organization of these microcircuits.
Our primary contributions are as follows:
  1. 1.
    We introduce a Variational Autoencoder (VAE)-based generative model specifically designed to learn a compressed latent representation of microcircuit topology from this mouse visual cortex data.
  2. 2.
    Crucially, we demonstrate that specific aspects of this learned latent space show strong, understandable relationships with key structural properties of the microcircuits, such as how densely connected they are or how they form clusters. This means we can find meaningful ways to describe the core variations in circuit structure.
  3. 3.
    Building on this, we propose a method for the controlled generation of microcircuits. By carefully adjusting these meaningful aspects in the latent space, our method can create new, artificial network structures that have specific desired characteristics.
We name our model NeuroStructGen, a Neural Structure Generator. To our knowledge, this combined approach of learning a compact representation of biological connectomes and then using its understandable features for precise, controllable generation of new circuits has not been done before. This work offers a new way to explore how changes in basic organizational rules might affect brain circuit structure. Furthermore, understanding the brain’s structure could eventually help in designing more efficient and capable artificial neural networks.
2 Problem Definition
The core motivation of this work stems from the “genetic bottleneck” hypothesis: the vast complexity of the brain’s connectome cannot be explicitly encoded by the genome, suggesting the existence of a compact, low-dimensional generative code. Our overarching goal is to develop a computational framework to learn, interpret, and control such generative representations of brain microcircuit topology. We define two primary objectives:
2.1 Learning Compact Generative Representations of Microcircuit Topology
Given a dataset of $N$ biological microcircuit graphs, denoted as $\mathbf{G}_{\text{data}}=\{\mathcal{G}^{(1)},\mathcal{G}^{(2)},\dots,\mathcal{G}^{(N)}\}$, the first objective is to learn a low-dimensional latent representation $\mathbf{z}^{(i)}\in\mathbb{R}^{d_{z}}$ for each microcircuit $\mathcal{G}^{(i)}$. Each microcircuit is represented as a graph $\mathcal{G}^{(i)}=(\mathbf{X}^{(i)},\mathbf{A}^{(i)})$, where $\mathbf{X}^{(i)}\in\mathbb{R}^{n_{i}\times d_{v}}$ is a matrix of node features for its $n_{i}$ neurons (with $d_{v}$ feature dimensionality), and $\mathbf{A}^{(i)}\in\{0,1\}^{n_{i}\times n_{i}}$ is its adjacency matrix indicating synaptic connections. In this work, since the number of nodes varies across the graph data, we pad each adjacency matrix to 100×100. For consistent processing, a canonical node ordering $\pi$ is assumed for inputs to certain model components, thus we may refer to an ordered graph as $\mathcal{G}_{\pi}^{(i)}$.
This learned latent representation $\mathbf{z}^{(i)}$ must be generative. That is, we aim to learn a probabilistic model $p_{\theta}(\mathcal{G}_{\pi}|\mathbf{z})$ capable of generating realistic microcircuit graphs from these latent codes. The quality of this representation will be assessed by its ability to:
  1. 1.
    Faithfully reconstruct observed graph structures and their topological properties from their latent codes.
  2. 2.
    Generate novel, diverse graphs that capture the statistical characteristics of the biological training data.
2.2 Controllable Generation of Microcircuits with Target Properties
Building upon the learned latent space $\mathcal{Z}$ (the space of all $\mathbf{z}$) and the generative model $p_{\theta}(\mathcal{G}_{\pi}|\mathbf{z})$ from Section 2.1, our second objective is to enable the controlled generation of novel microcircuits.
Specifically, given a target structural, dynamical, or functional property $\mathcal{T}$ (e.g., a specific mean degree, clustering coefficient, or level of assortativity), and a desired value $t_{\text{target}}$ for this property, the goal is to synthesize new connectome graphs $\mathcal{G}_{\pi,\text{new}}$. These generated graphs should:
  1. 1.
    Optimally satisfy the specified constraint, meaning the property $\mathcal{T}$ for $\mathcal{G}_{\pi,\text{new}}$ should be close to $t_{\text{target}}$.
  2. 2.
    Preserve general topological and dynamical characteristics inherent to biological neural microcircuits, ensuring they remain plausible.
Formally, this requires effectively sampling from, or being guided by, a conditional probability distribution $p(\mathbf{z}|\mathcal{T},t_{\text{target}})$ in the latent space. Latent vectors $\mathbf{z}_{\text{new}}$ drawn from this distribution are then decoded using $p_{\theta}(\mathcal{G}_{\pi}|\mathbf{z}_{\text{new}})$ to produce the desired microcircuits.
3 Related Work
Analyzing the structural organization of brain connectomes is fundamental to understanding the intricate relationship between brain structure and function ref1 ; ref2 . A number of methods have been used to address this link, including statistical models statistical_models , communication models communication_model1 , and biophysical models biophysical_models1 ; biophysical_models2 . By modeling brain networks as graphs, researchers employ graph-theoretic approaches to reveal key topological properties that underlie efficient communication and cognitive processes ref3 ; ref4 . Understanding how these structural features relate to functional dynamics is a central goal in connectomics.
To investigate the principles governing connectome formation and organization, generative models have been developed. Early approaches typically relied on a small set of predefined wiring rules or biological principles, such as cost-efficiency or growth mechanisms, to replicate observed network features in silico Generativemodelsofthehumanconnectome ; Modellingthedevelopmentofcorticalsystemsnetworks ; ASimpleGenerativeModel . While successful in capturing certain global properties, these rule-based methods often lack flexibility and depend on manually specified or constrained generative factors, limiting their ability to explore the full complexity of biological variability.
More recently, deep generative models, such as Variational Autoencoders (VAEs), have emerged as powerful tools for analyzing and synthesizing complex data like brain networks ref5 ; ref6 . These models are particularly well-suited for learning a low-dimensional latent representation of the network data, effectively performing dimensionality reduction. Furthermore, by manipulating or sampling from this learned latent space, these models enable the controlled synthesis of novel, biologically plausible network configurations, offering new avenues for exploring connectome variability and its functional implications ref7 ; ref8 .
4 Method
4.1 MICrONS Dataset and Preprocess
We adopt the IARPA MICrONS dataset microns , which encompasses a 1.4 × 0.87 × 0.84 mm volume of cortex from a P87 mouse microns , containing neurons within the area and their interconnections. Given the cortex’s division into six distinct layers, each associated with a specific level of information processing layers1 ; layers2 ; layer3 , our study concentrates on cortical microcircuits organized as vertically oriented functional columns that traverse all laminar layers circuit1 ; circuit2 . To model these microcircuits, we extracted cylindrical subvolumes oriented perpendicularly to the cortical layers. Each cylindrical unit comprises 80–100 neurons, preserving intra-column connectivity while excluding connections extending beyond the columnar boundary.
Each columnar circuit is treated as a binary directed graph. To maintain the relative positional relationship between neurons, we define an ordering rule $\pi$ based on their y-coordinates, representing their positions within the cortical hierarchy. This consistent sorting rule also facilitates the comparison between the original data and generated results during the training and evaluation of the generative model. The relative positions of neurons are encoded by one-hot vectors indicating their indices according to the defined order $\pi$. Neuron type and synapse weight information are temporarily disregarded, as this work focuses solely on the topological structure.
4.2 Connectome Graph Variational Autoencoder
As the goal is to find latent representations similar to “genetic bottleneck”, we need a generative model with “information bottleneck”. Thus, adopting a variational autoencoder (VAE) VAE on the connectome data can be formulated as follows: Assume a set of training connectome graphs $\mathbf{G}=\{\mathcal{G}_{\pi}^{(i)}\}$ is generated from the distribution of a set of unobserved latent representation $\textbf{z}=(z_{1},...,z_{m})$, where $z^{(i)}\sim p_{\theta^{*}}(z)$ and training data is sampled from true conditional $p_{\theta^{*}}(\mathcal{G}_{\pi}|z^{(i)})$. We denote the data generation process as $p_{\theta}(\mathcal{G}_{\pi}|\mathbf{z})$. Following the standard VAE setting VAE , we approximate the intractable posterior by $q_{\phi}(\mathbf{z}|\mathcal{G}_{\pi})\approx p_{\theta}(\mathbf{z}|\mathcal{G}_{\pi})$ and minimize the evidence lower bound on the marginal likelihood of graph $\mathcal{G}^{(i)}$:
$\log p_{\theta}(\mathcal{G}_{\pi}^{(i)})\geq\mathcal{L}(\phi,\theta;\mathcal{G}_{\pi}^{(i)})=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathcal{G}_{\pi}^{(i)})}\left[\log p_{\theta}(\mathcal{G}_{\pi}^{(i)}|\mathbf{z})\right]-\beta\mathrm{KL}\left[q_{\phi}(\mathbf{z}|\mathcal{G}_{\pi}^{(i)})||p_{\theta}(\mathbf{z})\right],$ (1)
where $\beta$ is a hyperparameter to balance the reconstruction loss and KL-divergence loss during the training process betaVAE .
The proposed VAE model for connectome graphs comprises four main components. First, a node feature encoder employs a multi-layer, multi-head Graph Attention Network (GAT) GAT ; attentionisallyouneed to compute an embedding vector for each node. Second, a graph global encoder, which is a transformer encoder augmented with a special token BERT ; pigvae ; attentionisallyouneed , extracts an embedding for the entire sequence of node embeddings. This extracted sequence embedding serves as the graph-level embedding, upon which a 32-dimensional mean and variance are computed. The reparameterization of a standard VAE is applied to this 32D latent embedding representing the whole graph. Third, a node feature decoder, which is a transformer decoder, takes the graph-level embedding as input to reconstruct node features necessary for subsequent edge prediction. Finally, the edge predictor utilizes these decoded node features to predict the edges of the graph.
The overall architecture of the VAE model is illustrated in Figure 1. The detailed encoder structure is depicted in Appendix Figure 8, and the detailed decoder structure is depicted in Appendix Figure 9.
Figure 1: Overall model structure.
Node Feature Encoder
The node feature encoder, utilizing a three-layer multi-head GAT network, transforms 100-dimensional one-hot node representations into 32-dimensional node embeddings. See Appendix 7.2 for GAT architecture details.
Graph Global Encoder
Treating the y-ordered nodes as a sequence (analogous to words in a sentence), the graph global encoder transforms node embeddings into a fixed-size latent representation. Following BERT ; pigvae , a dummy node $v_{0}$ is prepended to the sequence to serve as a global embedding. The encoder applies rotational positional encoding (RoPE) rope and several transformer encoder layers to this augmented sequence. The final embedding of $v_{0}$ is taken as the global graph representation. An MLP is then applied to compute the 32-dimensional mean and variance for sampling the 32-dimensional latent vector $\mathbf{z}$, similar to standard VAEs.
Node Feature Decoder
The Node Feature Decoder reconstructs individual node embeddings from the global graph embedding using several transformer decoder layers. The input is the global graph embedding, augmented with rotational positional encoding (RoPE) rope . The global graph embedding also serves as memory for the decoder’s cross-attention, leveraging this global context during node feature reconstruction.
Edge Predictor
The edge predictor is a cross-node interaction layer. It takes the node feature decoder output $h\in\mathbb{R}^{n\times d}$, where $n=100$ is the maximum number of nodes and $d$ is the embedding dimension. Edges are predicted using the dot product of embeddings transformed by two distinct linear layers with activation.
$\mathbf{A}_{\text{pred}}=\sigma(\text{LeakyReLU}(\mathbf{h}\mathbf{W}_{1})(\text{LeakyReLU}(\mathbf{W}_{2}\mathbf{h})^{\top})),$ (2)
The output $\mathbf{A}_{\text{pred}}$ provides a probabilistic adjacency matrix where each entry, a floating-point number between 0 and 1, denotes the likelihood of an edge. We then perform a Bernoulli sampling process on each entry using this probability to generate a binary adjacency matrix.
4.3 Controllable Connectome Generation by Sampling from the Latent Space
Investigating the intricate interplay between structure, dynamics, and function in brain neural microcircuits necessitates the capability to generate novel networks in a controlled fashion. Specifically, given a target network property $\mathcal{T}$ (which can be structural, dynamical, or functional), our objective is to synthesize novel connectomes $\mathcal{G}_{\pi}$ that optimally satisfy the specified constraint $\mathcal{T}$ while preserving general topological and dynamical characteristics inherent to biological neural microcircuits. Formally, the goal is to generate new samples from the conditional probability distribution $p(\mathcal{G}_{\pi}|\mathcal{T})$.
Given the computational cost associated with large-scale graph generation via brute-force sampling, we propose leveraging latent space properties to inform our sampling strategy. By employing geometric characteristics of the latent manifold as a sampling heuristic, this approach significantly reduces computational overhead compared to the brute-force generate-then-filter paradigm. Specifically, we aim to find the probability distribution of the latent vector $\mathbf{z}$ conditioned on the target property $\mathcal{T}$, denoted as $p(\mathbf{z}|\mathcal{T})$, which can be represented as an energy model:
$p(\mathbf{z}|\mathcal{T})=\frac{1}{Z}p(\mathbf{z})^{{1}/{\tau}}\exp(\lambda S(\mathcal{T},\mathbf{z})),$ (3)
[…]
When $\lambda\to\infty$, the energy model degenerate into a strictly constrained version (by $\mathcal{T}$). In this case we have:
$p(\mathbf{z}|\mathcal{T})=\frac{p(\mathbf{z})^{{1}/{\tau}}}{p(\Omega_{\mathcal{T}})}\cdot\mathbb{I}(\mathbf{z}\in\Omega_{\mathcal{T}}),$ (4)
[…]
5 Experiments
5.1 Evaluation Metrics
We selected several graph theory metrics (descriptors) commonly employed in the analysis of connectomes brain_graph_metrics , including mean degree, efficiency, transitivity, clustering coefficient, modularity, and assortativity. The detailed definitions of these metrics are provided in the Appendix 7.3.
5.2 Microcircuit Reconstruction and Generation Results
Examples of original and reconstructed graphs obtained through the VAE are provided in the Appendix 7.4. To validate the model’s generative capabilities, we evaluated several graph metrics on the generated samples. We employed the maximum mean discrepancy (MMD) MMD to compare the distributions of these graph statistics between an equal number of generated and test graphs. Following the methodology in you2018b_mmd , we specifically measured the distributions of degree, clustering coefficient and spectrum. For MMD computation, we utilized both the Gaussian Earth Mover’s Distance (EMD) kernel and the total variation (TV) distance liao2019 . Given the absence of existing generative models specifically designed for connectome graphs, we selected three alternative models commonly used for molecule generation or general graph generation GDSS ; DisCo ; EDGE as baselines. The following Table 1 reports the MMD values for different graph metrics evaluated across these models and our proposed approach. Figure 2 compares graphs generated by our proposed model and other models with the original data (column Ori.).
Model Deg. Clus. Coef. Spec.
EMD $\downarrow$ TV $\downarrow$ EMD $\downarrow$ TV $\downarrow$ EMD $\downarrow$ TV $\downarrow$
GDSS 1.138 0.493 1.381 0.959 0.325 0.515
DisCo 1.047 0.304 1.315 0.550 0.094 0.334
EDGE 0.660 0.041 0.343 1.016 0.972 0.111
NeuroStructGen (Ours) $\boldsymbol{0.164}$ $\boldsymbol{0.028}$ $\boldsymbol{0.154}$ $\boldsymbol{0.021}$ $\boldsymbol{0.047}$ $\boldsymbol{0.007}$
Table 1: Comparison with Baseline Models
(a) Ori.
(b) Ori.
(c) Ours
(d) Ours
(e) GDSS
(f) GDSS
(g) EDGE
(h) EDGE
(i) DisCo
(j) DisCo
Figure 2: Generated result comparison.
5.3 Latent Representation Analysis
5.3.1 SHAP Analysis
Using SHAP analysis shap to interpret feature contributions, we analyzed the relationship between the 32 latent dimensions and generated graph properties. Figure 3 illustrates the SHAP results for the clustering coefficient, showing that a few dimensions largely control this metric. Similar analyses were performed for each graph property (see Appendix 7.5).
Figure 3: SHAP analysis for clustering coefficient.
5.3.2 Discovering Key Latent Directions Affecting Generation Metrics
To visualize the main axes of topological structure variation across the microcircuits, we applied non-linear dimensionality reduction using t-SNE and identified topological features exhibiting gradients in the resulting embedding space (Figure 4). For each graph descriptor, we then aimed to find a consistent direction in the latent space that correlates with changes in its value. To achieve this, we divided the 500 test graphs into 20 value bins, each containing 5% of the data, and trained a linear regression model to predict the bin index from the 32-dimensional latent vector. The constant gradient direction of the linear regression model is advantageous for identifying this direction linear_reg , and further details of this method are in Appendix 7.6. We plot the $\text{R}^{2}$ scores of these linear regression models in Figure 5(a).
Since there might be some inherent correlation between the metrics, which results in similar or divergent gradient directions for different metrics, we computed the correlation between each pair of metrics directly from the test dataset, as shown in Figure 5(b). We also computed the cosine similarity between each pair of gradient directions, as shown in Figure 5(c). The relative magnitude of the cosine similarity between the correlation values and the gradient directions is largely consistent, further confirming that the gradient direction does indeed correspond to specific metrics. Further details regarding the correlations between different metrics are provided in Appendix7.8.
To validate whether these identified directions truly reflect changes in the corresponding metrics, we performed a traversal experiment. Starting from the mean latent vector of the test set, we moved along each metric’s identified gradient direction and decoded the resulting latent vectors to observe the change in the metric’s value. The details are shown in Appendix 7.7.
Refer to caption(a) Mean degreeRefer to caption(b) EfficiencyRefer to caption(c) TransitivityRefer to caption(d) AssortativityRefer to caption(e) Clus. Coef.Refer to caption(f) Modularity
Figure 4: t-SNE visualizations of vector embeddings reveal axes of variation in connectome topology. The darkness of each point’s color corresponds to the magnitude of its associated bin index for the specific metric being considered. To further aid visualization, the red line overlaid on the t-SNE plot represents the direction of metric variation, obtained by fitting an auxiliary 2D linear regression model to the 2D t-SNE embeddings. This 2D regression is solely for visual convenience and does not replace the 32-dimensional analysis described in the main text.
Refer to caption(a) $\text{R}^{2}$ scoresRefer to caption(b) CorrelationsRefer to caption(c) Gradient distance
Figure 5: (a) $\text{R}^{2}$ scores of the linear regression models predicting each of the six graph metrics from the 32-dimensional latent embeddings, indicating the predictive strength of the latent space. (b) Spearman’s rank correlation coefficient matrix between the six graph metric descriptors, calculated directly from the test dataset. (c) Cosine distance matrix between the gradient directions in the 32-dimensional latent space for each of the six graph metrics, illustrating the similarity in how different metrics are encoded in the latent space.
5.4 A Pipeline for Controllable Microcircuit Generation With MCMC Sampling
For the investigation of structure-function relationships in brain networks, the ability to generate network samples that closely match specific target descriptor values would be invaluable. Here, we propose a general pipeline to achieve this goal, leveraging our latent VAE connectome generator in conjunction with Markov Chain Monte Carlo (MCMC) sampling MCMC .
[…]
As direct integration to compute $p(\Omega_{\mathcal{T}})$ is not feasible, we utilize Markov Chain Monte Carlo (MCMC) MCMC to sample from the target conditional probability distribution despite the unknown denominator. We define a log-target-density function (LTD) to evaluate the likelihood of a proposed latent vector $\mathbf{z}$:
$\text{LTD}(\textbf{z})=\displaystyle\frac{1}{\tau}\text{log}p(\textbf{z}),$ (5)
Thus, the target conditional distribution for should be proportional to $\text{exp}(\text{LTD}(\mathbf{z}))\cdot\mathbb{I}(\mathbf{z}\in\Omega_{\mathcal{T}})$. We utilize the Metropolis-Hastings acceptance rule to generate a sequence of $N$ sampled latent vectors. The detailed steps of the sampling algorithm are provided in the Appendix 7.9.
To demonstrate the ability to generate graphs satisfying specific properties, we conducted experiments by specifying target percentile values ranging from 0% to 100%. We set $N=500$ in the experiment. The mean values of the graph metrics for the 500 generated graphs at different target percentiles are shown in Figure 6. Generated graph metric values that show a roughly monotonically increasing trend with the target bin index demonstrates the effectiveness of this method. Furthermore, examples of the generated graphs for different target metrics and target values are presented in the Appendix7.10.
Figure 6: Metrics of generated graphs when setting different targets.
Figure 7 presents examples of controlled graph generation where the mean degree is the targeted property, showing samples generated by setting different target percentile ranges for this metric. Detailed results for controlling other graph properties are provided in the Appendix 7.10.
(a) target=0%
(b) 20%
(c) 40%
(d) 60%
(e) 80%
(f) 100%
Figure 7: Generated graph examples targeting different mean degree percentile ranges.
6 Conclusions and Discussions
In this work, we introduced a VAE-based generative model that successfully learns a compact and interpretable latent representation of mouse cortical microcircuit topology from detailed connectome data. We demonstrated that specific, interpretable directions within this latent space correspond to meaningful network properties, enabling the controlled synthesis of novel microcircuits with desired structural features via latent space navigation and an MCMC-based pipeline. This approach offers a powerful new tool to investigate the underlying design principles of neural circuits, explore structure-function relationships, and potentially inform the development of more advanced artificial neural networks by elucidating how complex structures might arise from low-dimensional generative rules.
Limitations and Future Work
Despite promising results, this work has several limitations. Firstly, our choice of a Variational Autoencoder was deliberate, driven by the primary objective of learning an explicit, interpretable, and low-dimensional latent space. This “information bottleneck” is crucial for uncovering the compact generative blueprint hypothesized to underlie brain development and for enabling controlled synthesis. While other modern generative models, such as diffusion models, demonstrate powerful sample generation capabilities, their latent spaces are not always as directly optimized for, or as easily amenable to, the extraction of such a compressed and interpretable code as a VAE’s. Future work could, however, investigate adaptations of these models or hybrid approaches to achieve similar goals. Secondly, our model currently focuses on binary topological structure, neglecting crucial biological details such as neuron types, synaptic weights, and activity dynamics. Incorporating these features is a key direction for future work to enhance biological realism. Thirdly, the study is based on microcircuits from mouse visual cortex; the learned representations and generative rules may need adaptation and validation for other brain regions, larger-scale connectomes, or different species. The current fixed-size input representation (100x100 padding) also poses challenges for direct application to circuits of highly variable sizes without architectural modifications. Furthermore, while we identified interpretable linear directions in the latent space, exploring more complex, non-linear relationships and the full extent of encoded biological constraints warrants further investigation. Finally, validating the functional viability of generated circuits, for instance through in-silico simulations, is an important next step to bridge the gap between structural generation and functional understanding.